OpenStudy (anonymous):

How can I write a good analysis on this problem y=e^x because my calculus teacher told me to use excel and put 100 different numbers for x from 0 to 5. So I just increases each number by 0.05...But I dont know how to write a analysis on this problem! It has to be a page long! He said I could write about how to make it better but how would I do that? Please help!!!

OpenStudy (unklerhaukus):

y=e^x is not a problem, y=e^x is an equation what are you trying to doing with this equation ?

OpenStudy (anonymous):

sorry about that....My calculus teacher wants me to take that equation and put in 100 different numbers for x...From 0-5 he said....So I took 100 into 5 and got a increase of 0.05 each time. He wants me to use excel and write a analysis on how to make the answer better and other stuff...But I dont know how! :/

OpenStudy (unklerhaukus):

the answer to what ? what are you asked to find ?

OpenStudy (anonymous):

I meant equation...How can I make the equation better?

OpenStudy (unklerhaukus):

?